# The genealogy of nearly critical branching processes in varying environment

### v3

@@ 2.2 Examples
Let us give some examples illustrating the conditions (6) and (7) which fit the class of models which we are considering here.
@@
    with $\alpha\in\mathbb{R}$ and $\sigma^{2}>0$. We then have
@@
    in the Skorohod topology, and the variance process (7) converges to $\sigma^{2}(t)=\sigma^{2}t$.
@@
    with $\alpha\in\mathbb{R}$, $\sigma^{2}>0$ and some zero mean random variable $\zeta$ with finite second moment. Then
@@
    in law for the Skorohod topology, where $(B_{s},s\geq 0)$ is a standard Brownian motion and $\alpha^{\prime},\rho$ are some constants depending on the moments of $\zeta$.
@@
    The setting of the second example is easily adapted to cases where the corresponding random walk converges to a general Lévy process instead.
@@
    Let $p\in(0,1)$. Consider a varying environment where
@@
    Elsewhere, suppose that the environment is constant with mean $1$ and variance $\beta$. Then the variance process converges to the linear function $\beta t$ with a jump of size $1$ at $1/2$. Note that (8) is fulfilled, whereas (8) is only fulfilled for $p<1/2$. By the same token, one can extend the construction to generate limiting varying environments where $\sigma^{2}$ is a subordinator and $X$ an arbitrary independent Lévy process.
@@ 2.3 Main results
Let $(Z_{n},n\geq 0)$ be a sequence of BPVEs. For $t>0$ such that $t$ is a continuity point of $\sigma^{2}$ and $\sigma^{2}$ strictly increasing at $t$, then as $N\to\infty$ under the near critical assumptions (6), (7) and (8) we have
$\mathbb{P}(Z_{tN}>0)\sim\frac{2}{\kappa_{N}}\frac{1}{\int_{[0,t]}e^{-X_{s}}\sigma^{2}(ds)}.$ (14)
We note that in the case of critical or nearly critical Galton–Watson processes, Theorem 1 reduces to well known results, see [37] and [47]. The proof of Theorem 1 will be given in Section 6.
@@
which corresponds to the time to the most-recent common ancestor of $u$ and $v$. Consider the empirical measure
@@
Then $[T_{n},d_{n},\lambda_{n}]$ is a random metric measure space that encodes the genealogy of generation $n$ of the BPVE. Formally, we can define a topology on the space of metric measure spaces, the Gromov–Hausdorff–Prohorov topology [1], and $[T_{n},d_{n},\lambda_{n}]$ is a random element of this space. The definition of this topology is recalled in Section 3.
In order to state our main result we briefly introduce here the scaling limit of the latter metric measure space, and name it the environmental coalescent point process (CPP). We construct it as a time-change of the more standard Brownian CPP [48]. A more direct construction is provided in Section 4. Consider a Poisson point process $P$ on $[0,\infty)\times(0,1]$ with intensity measure $ds\otimes\tfrac{1}{x^{2}}dx$. Define a distance on $\mathbb{R}_{+}$ as
@@
Consider an independent standard exponential random variable $Z_{B}$. The Brownian CPP is the random metric measure space $[(0,Z_{B}),d_{B},\operatorname{Leb}]$, where $\operatorname{Leb}$ is the Lebesgue measure on the interval $(0,Z_{B})$. It corresponds to the universal scaling limit of genealogies of critical branching processes with finite variance and is illustrated in Figure 1.
Figure 1: Simulation of a Brownian CPP. Left, a vertical line of height $x$ at location $s$ represents an atom $(s,x)$ of $P$. Right, the tree corresponding to the left CPP. The distance $d_{B}$ is the tree distance on the leaves.
@@
$[(0,Z_{e}),d_{\nu_{e}},\operatorname{Leb}]\coloneqq[(0,Z_{B}),F^{-1}\circ d_{B},\bar{\rho}_{t}\operatorname{Leb}].$
@@
Suppose that (6), (7) and (8) hold and fix $t>0$ that fulfils the same assumption as in Theorem 1. Conditional on survival at time $tN$ the following holds in distribution for the Gromov–Hausdorff–Prohorov topology
@@
with $[(0,Z_{e}),d_{\nu_{e}},\operatorname{Leb}]$ the environmental CPP.
This result is proved in Section 7. We want to point at the following fact. The limiting genealogy, the environmental CPP, is obtained by time-changing the Brownian CPP according to the function $F$ which might be discontinuous (as in the last example of Section 2.2). As a consequence, although the Brownian CPP is a binary tree, the environmental CPP might have (simultaneous) multiple mergers. These multiple mergers correspond to an accumulation of binary branch points over a time-scale shorter than $N$, due to a region of highly variable environments. This phenomenon is illustrated in Figure 2.
Figure 2: Simulation of the genealogy of a BPVE until generation $N=500$. All offspring distributions are negative binomials with mean $1+\alpha/N$, with $\alpha=1$. Outside of the shaded region, the variance is $\sigma^{2}=2$. In the shaded region, the variance is increased in such a way that the sum of the variances in that region is $cN$, with $c=2$.
@@ 2.4 Consequences of Theorem 2
The convergence of the population in the GHP topology might seem daunting at first sight to the reader not familiar with random metric measure spaces. We show in this section how the GHP convergence in Theorem 2 condenses the limit of several genealogical quantities of interest. This requires some preliminary notation for these quantities and their limits.
For each $i\leq tN$, let $Z_{i,tN}$ be the number of individuals at generation $i$ having descendants at generation $tN$ in the BPV individuals chosen uniformly at random from $T_{tN}$ (the set of individuals at generation $tN$) and recall that $d_{tN}(U_{i},U_{j})$ denotes the time to the most-recent common ancestor of $U_{i}$ and $U_{j}$.
For $k\in\mathbb{N}$ let $(H_{1},\dots,H_{k-1})$ be a sequence of independent random variables with cumulative distribution function $F$ defined in (16). We define the array $(H_{i,j};i,j\leq k)$ via
$\forall i\leq j,\quad H_{i,j}=H_{j,i}=\max\{H_{i},\dots,H_{j-1}\},$ (18)
with the convention that $\max\varnothing=0$. Additionally, for $\theta\geq 0$ we introduce $H^{\theta}$ with
@@
Lastly, we introduce the array $(H^{\theta}_{i,j};i,j\leq k)$ in the same manner as in (18). We note that (19) in fact defines a valid probability distribution for any $\theta>-1$. However, these values of $\theta$ are not relevant in our case. Some motivation for this distribution in connection with coalescent point processes is provided in Section 4.2.
@@
Let $t>0$ fulfil the same assumptions as in Theorem 1. The following limits hold in distribution as $N\to\infty$, conditional on $Z_{Nt}>0$.
@@
    where $\mathscr{E}(x)$ is an exponentially distributed random variable with mean $x>0$.
@@
    in the sense of finite-dimensional distribution at continuity points of $\sigma^{2}$, and where $(Y(t),t\geq 0)$ is a Yule process. If $\sigma^{2}$ is continuous, the convergence holds in the Skorohod sense.
@@
    $\big{(}\tfrac{d_{tN}(U_{i},U_{j})}{N}\big{)}_{i,j}\to(H_{\sigma_{i},\sigma_{j}})_{i,j},$
    where $\sigma$ is a uniform permutation of $\{1,\dots,k\}$ independent of $(H_{i,j})_{i,j}$, and the law of the latter matrix is given by
[…]
    where $\big{(}(H^{\theta}_{i,j})_{i,j}\big{)}$ has distribution (19) and $\phi\colon\mathbb{R}_{+}^{k\times k}\to\mathbb{R}_{+}$ is continuous bounded.
@@
Finally, let us illustrate our result by specifying the previous corollary to our first example, a nearly critical Galton–Watson process with mean $1+\alpha/N+o(1/N)$ and variance $\sigma^{2}+o(1)$. Theorem 1 becomes
@@
Note that the case $\alpha=0$ corresponds to the classical Kolmogorov asymptotics [37], whereas the case $\alpha\neq 0$ can be found in [47]. Furthermore, the first point of Corollary 2.3 yields that, conditional on $Z_{tN}>0$,
@@
These are well-known asymptotics for nearly critical branching processes, see for instance [47]. In turn, point (iii) of Corollary 2.3 and a small calculation prove that the joint limiting distribution of the split times $s_{1},\dots,s_{k-1}\in[0,t]^{k-1}$ of the subtree spanned by $(U_{1},\dots,U_{k})$ has density $f$ given by
@@
which is exactly the limit in [26, Theorem 3] for $\alpha=0$. Finally, point (ii) shows that the reduced process is asymptotically distributed as the time-changed Yule process,
@@
The main results in [33] and [24] are respectively that, after making a time-rescaling $u=G^{(N)}(s)$, the reduced process and the law of a $k$ sample from a BPVE converge to the corresponding expressions for the Brownian CPP. Since $G^{(N)}(s)\to G(s)\coloneqq\frac{\bar{\rho}_{s}}{\bar{\rho}_{t}}$, (see Lemma A.4), reverting the limiting time-change, that is taking $s=F^{-1}(u)$, yields the environmental CPP (see Proposition 4.2). Thus points (ii) and (iii) of Corollary 2.3 are consistent with the results obtained in [33] and [24]. Note again that we work under different settings, see Remark 2.2.
@@ 3 Random metric measure spaces and Gromov topologies
The bulk of our work is dedicated to proving convergence in the Gromov-weak topology which is introduced in Section 3.1.1. In this topology, convergence in distribution is strongly connected to convergence of moments of the metric space, which are computed out of samples from the metric space. These moments can be computed efficiently for branching processes using the many-to-few formula of Section 5.2. Applying this method of moments requires the reproduction laws to have moments of any order, whereas our convergence result only requires a finite moment of order two. To overcome this difficulty, we use a truncation procedure similar to that in [26]. Section 3.1.2 contains some perturbation results that are used to prove that the truncated version of the BPVE remains close in the Gromov-weak topology to the original BPVE.
@@ 3.1.1 The Gromov-weak topology
A metric measure space is a triple $[X,d,\nu]$ such that $(X,d)$ is a complete separable metric space and $\nu$ is a finite measure on the corresponding Borel $\sigma$-field. Given some $k\geq 1$ and a bounded continuous map $\phi\colon[0,\infty)^{k\times k}\to\mathbb{R}$, we define a functional $\Phi$ on the space of metric measure spaces, called a polynomial, as
@@
The Gromov-weak topology is defined as the topology induced by the polynomials, that is, the smallest topology making all polynomials continuous. The Gromov-weak topology allows us to define a random metric measure space as a random variable with values in metric measure spaces, endowed with the Gromov-weak topology and the corresponding Borel $\sigma$-field. The moment of a random metric measure space associated to $\Phi$ is the expectation of the corresponding polynomial.
@@
Then $[T_{n},d_{n},\lambda_{n}]$ is the random metric measure space encoding the genealogy of generation $n$ of the BPVE.
@@
for some limiting space $[X,d,\nu]$. Then if the limiting space further fulfils that
@@
the sequence $([X_{n},d_{n},\nu_{n}],n\geq 1)$ converges to $[X,d,\nu]$ in distribution for the Gromov-weak topology.
@@
Let us compare the previous result with the usual method of moments for real random variables. Let $(X_{n})$ be a sequence of real r.v.s and assume that for every $k\in\mathbb{N}$ there exists some $m_{k}$ such that
@@
Subject to an additional technical condition, the method of moments leads to two key outcomes: (1) the $m_{k}$’s coincide with the moments of an underlying random variable $X_{\infty}$, and (2) the sequence $(X_{n})$ converges in distribution to $X_{\infty}$. This stands in contrast to the prior result where (1) is not guaranteed; specifically, the limiting moments are required to coincide with the moments of a limiting metric space. Notably, Proposition 3.1 can be enhanced by relaxing this requirement, as detailed in [15], Section 4. This improvement hinges on extending the Gromov-weak topology to non-separable ultrametric spaces through a de Finetti representation for exchangeable coalescents, as obtained in [14].

### v4

@@ 2.2 Examples
We now present some examples illustrating the conditions (6) and (7), all of which fit the class of models we are considering here.
@@
    with $\alpha\in\mathbb{R}$ and $\sigma^{2}>0$. Then
@@
    in the Skorohod topology as $N\to\infty$, and the variance process (7) converges to $\sigma^{2}(t)=\sigma^{2}t$.
@@
    with $\alpha\in\mathbb{R}$, $\sigma^{2}>0$ and some mean-zero random variable $\zeta$ with finite second moment. Then
@@
    in law in the Skorohod topology as $N\to\infty$, where $(B_{s},s\geq 0)$ is a standard Brownian motion and $\alpha^{\prime},\rho$ are some constants depending on the moments of $\zeta$.
@@
    The setting of the second example is easily adapted to cases in which the corresponding random walk converges to a general Lévy process instead.
@@
    Let $p\in(0,1)$. Consider a varying environment such that
@@
    Elsewhere, suppose that the environment is constant with mean $1$ and variance $\beta$. Then the variance process converges to the linear function $\beta t$ with a jump of size $1$ at $1/2$. Note that (8) is fulfilled, whereas (8) holds only for $p<1/2$. By the same token, one can extend the construction to generate limiting varying environments for which $\sigma^{2}$ is a subordinator and $X$ an arbitrary independent Lévy process.
@@ 2.3 Main results
Let $(Z_{n},n\geq 0)$ be a sequence of BPVEs satisfying (6), (7), and (8). For any $t>0$ such that $\sigma^{2}$ is continuous and strictly increasing at $t$,
$\mathbb{P}(Z_{tN}>0)\sim\frac{2}{\kappa_{N}}\frac{1}{\int_{[0,t]}e^{-X_{s}}\sigma^{2}(ds)},\qquad\text{as $N\to\infty$.}$ (14)
We note that in the case of critical or nearly critical Galton–Watson processes, Theorem 1 reduces to well-known results, see [37] and [47]. The proof of Theorem 1 will be given in Section 6.
@@
which is equal to the time to the most-recent common ancestor of $u$ and $v$. Consider the empirical measure
@@
Then $T_{n},d_{n},\lambda_{n}]$ is a random metric measure space that encodes the genealogy of generation $n$ of the BPVE. Formally, we can define a topology on the space of metric measure spaces, the Gromov–Hausdorff–Prohorov topology [1], and $T_{n},d_{n},\lambda_{n}]$ is a random element of this space. The definition of this topology is recalled in Section 3.
In order to state our main result we briefly introduce here the scaling limit of the latter metric measure space, and name it the environmental coalescent point process (CPP). We construct it as a time-change of the standard Brownian CPP [48]. A more direct construction is provided in Section 4. Consider a Poisson point process $P$ on $0,\infty)\times(0,1]$ with intensity measure $ds\otimes\tfrac{1}{x^{2}}dx$. Define a distance on $\mathbb{R}_{+}$ as
@@
Consider an independent standard exponential random variable $Z_{B}$. The Brownian CPP is the random metric measure space $(0,Z_{B}),d_{B},\operatorname{Leb}]$, where $\operatorname{Leb}$ is the Lebesgue measure on the interval $(0,Z_{B})$. It corresponds to the universal scaling limit of genealogies of critical branching processes with finite variance and is illustrated in Figure 1.
Figure 1: Simulation of a Brownian CPP. Left, a vertical line with height $x$ at horizontal location $s$ represents an atom $(s,x)$ of $P$. Right, the tree corresponding to the CPP on the left. The distance $d_{B}$ is the tree distance on the leaves.
@@
$(0,Z_{e}),d_{\nu_{e}},\operatorname{Leb}]\coloneqq[(0,Z_{B}),F^{-1}\circ d_{B},\bar{\rho}_{t}\operatorname{Leb}].$
@@
Suppose that (6), (7) and (8) hold and fix $t>0$ that fulfils the same assumption as in Theorem 1. Conditional on survival at time $tN$, the following convergence holds in distribution for the Gromov–Hausdorff–Prohorov topology
@@
with $(0,Z_{e}),d_{\nu_{e}},\operatorname{Leb}]$ the environmental CPP.
This result is proved in Section 7. We point out the following fact. The limiting genealogy, the environmental CPP, is obtained by time-changing the Brownian CPP according to the function $F$ which might be discontinuous (as in the last example of Section 2.2). As a consequence, although the Brownian CPP is a binary tree, the environmental CPP might have (simultaneous) multiple mergers. These multiple mergers correspond to an accumulation of binary branch points over a time-scale shorter than $N$, due to a region of highly variable environments. This phenomenon is illustrated numerically in Figure 2.
Figure 2: Simulation of the genealogy of a BPVE until generation $N=500$. All offspring distributions are negative binomials with mean $1+\alpha/N$, with $\alpha=1$. Outside of the shaded region, the variance is $\sigma^{2}=2$. In the shaded region, the variance is increased so that the sum of the variances in that region is cNcN, with $c=2$.
@@ 2.4 Consequences of Theorem 2
The convergence of the population in the GHP topology might seem daunting at first sight to a reader not familiar with random metric measure spaces. We show in this section how the GHP convergence in Theorem 2 condenses the limit of several genealogical quantities of interest. This requires some preliminary notation for these quantities and their limits.
For each $i\leq tN$, let $Z_{i,tN}$ be the number of individuals at generation $i$ having descendants at generation $tN$ in the BPVE. Recall that, if $(U_{i})_{i\leq k}$ are chosen uniformly from $T_{tN}$ (the set of individuals alive at generation $tN$), $d_{tN}(U_{i},U_{j})$ denotes the time to the most-recent common ancestor of $U_{i}$ and $U_{j}$. For $k\in\mathbb{N}$ and $\theta\geq 0$, let $(H^{\theta}_{1},\dots,H^{\theta}_{k-1})$ be independent random variables with cumulative distribution function
@@
We define an array $(H^{\theta}_{i,j};i,j\leq k)$ via
$\forall i\leq j,\quad H^{\theta}_{i,j}=H^{\theta}_{j,i}=\max\{H^{\theta}_{i},\dots,H^{\theta}_{j-1}\},$ (19)
with the convention that $\max\varnothing=0$. We note that (18) in fact defines a valid probability distribution for any $\theta>-1$. However, these values of $\theta$ in (1,0)(-1,0) are not relevant in our case. Some motivation for this distribution in connection with coalescent point processes is provided in Section 4.2.
@@
Let $t>0$ satisfy the same assumptions as in Theorem 1. Conditional on $Z_{Nt}>0$, the following limits hold in distribution as $N\to\infty$.
@@
    where $\mathscr{E}(x)$ is an exponential random variable with mean $x>0$.
@@
    in the sense of finite-dimensional distribution at continuity points of $\sigma^{2}$, where $(Y(t),t\geq 0)$ is a Yule process. If $\sigma^{2}$ is continuous, the convergence holds in the Skorohod sense.
@@
    $\big(\tfrac{d_{tN}(U_{i},U_{j})}{N}\big)_{i,j}\to(\widetilde{H}_{\sigma_{i},\sigma_{j}})_{i,j},$
    where $\sigma$ is a uniform permutation of $\{1,\dots,k\}$ and $(\widetilde{H}_{i,j})_{i,j}$ and independent array distributed as
    $\mathbb{E}\big[\phi\big((\widetilde{H}_{i,j})_{i,j}\big)\big]=k\int_{0}^{\infty}\frac{1}{(1+\theta)^{2}}\Big(\frac{\theta}{1+\theta}\Big)^{k-1}\mathbb{E}\big[\phi\big((H^{\theta}_{i,j})_{i,j}\big)\big]d\theta,$ (20)
    where $\big((H^{\theta}_{i,j})_{i,j}\big)$ is defined in (19) and $\phi\colon\mathbb{R}_{+}^{k\times k}\to\mathbb{R}_{+}$ is continuous bounded.
@@
Finally, we illustrate our result by specifying the previous corollary to our first example, a nearly critical Galton–Watson process with mean $1+\alpha/N+o(1/N)$ and variance $\sigma^{2}+o(1)$. Theorem 1 becomes
@@
Note that the case $\alpha=0$ corresponds to the classical Kolmogorov asymptotics [37], whereas the case $\alpha\neq 0$ can be found in [47]. Furthermore, point (i) of Corollary 2.3 yields that, conditional on $Z_{tN}>0$,
@@
These are well-known asymptotics for nearly critical branching processes, see for instance [47]. Point (iii) of Corollary 2.3 together with a small calculation shows that the joint limiting distribution of the split times $s_{1},\dots,s_{k-1}\in[0,t]^{k-1}$ of the subtree spanned by $(U_{1},\dots,U_{k})$ has density $f$ given by
@@
which is exactly the limit in [25, Theorem 3] for $\alpha=0$. Finally, point (ii) shows that the reduced process is asymptotically distributed as the time-changed Yule process,
@@
The main results in [33] and [26] show that after the time-rescaling $u=G^{(N)}(s)$, the reduced process and the law of a $k$ sample from a BPVE converge to the corresponding expressions for the Brownian CPP. Since $G^{(N)}(s)\to G(s)\coloneqq\frac{\bar{\rho}_{s}}{\bar{\rho}_{t}}$, (see Lemma A.4), reverting the limiting time-change, that is taking $s=F^{-1}(u)$, yields the environmental CPP (see Proposition 4.2). Thus points (ii) and (iii) of Corollary 2.3 are consistent with the results obtained in [33] and [26]. Note again that we work under different settings, see Remark 2.2.
@@ 3 Random metric measure spaces and Gromov topologies
The bulk of our work is dedicated to proving convergence in the Gromov-weak topology, which is introduced in Section 3.1.1. In this topology, convergence in distribution is strongly connected to convergence of moments of the metric space, which are computed out of samples from the metric space. These moments can be computed efficiently for branching processes using the many-to-few formula of Section 5.2. Applying this method of moments requires the reproduction laws to have moments of any order, whereas our convergence result only requires a finite moment of order two. To overcome this difficulty, we use a truncation procedure similar to that in [25]. Section 3.1.2 contains some perturbation results that are used to prove that the truncated version of the BPVE remains close in the Gromov-weak topology to the original BPVE.
@@ 3.1.1 The Gromov-weak topology
A metric measure space is a triple $X,d,\nu]$ such that $(X,d)$ is a complete separable metric space and $\nu$ is a finite measure on the corresponding Borel $\sigma$-field. Given some $k\geq 1$ and a bounded continuous map $\phi\colon[0,\infty)^{k\times k}\to\mathbb{R}$, we define a functional $\Phi$ on the space of metric measure spaces, called a polynomial, as
@@
The Gromov-weak topology is defined as the topology induced by the polynomials, that is, the smallest topology making all polynomials continuous. The Gromov-weak topology allows us to define a random metric measure space as a random variable with values in metric measure spaces, endowed with the Gromov-weak topology and the corresponding Borel $\sigma$-field. The moment of a random metric measure space corresponding to $\Phi$ is the expectation of the corresponding polynomial.
@@
Then $T_{n},d_{n},\lambda_{n}]$ is the random metric measure space encoding the genealogy of generation $n$ of the BPVE.
@@
for some limiting space $X,d,\nu]$. Then if the limiting space further fulfils that
@@
the sequence $([X_{n},d_{n},\nu_{n}],n\geq 1)$ converges to $X,d,\nu]$ in distribution for the Gromov-weak topology.
@@
Let us compare the previous result with the usual method of moments for real random variables. Let $(X_{n})_{n}$ be a sequence of real r.v.s and assume that for every $k\in\mathbb{N}$ there exists some $m_{k}$ such that
@@
Subject to an additional technical condition, the method of moments leads to two key outcomes: (1) the $m_{k}$’s coincide with the moments of an underlying random variable $X_{\infty}$, and (2) the sequence $(X_{n})_{n}$ converges in distribution to $X_{\infty}$. This is in contrast with the result above, for which (1) is not guaranteed; specifically, the limiting moments are required to coincide with the moments of a limiting metric space. Notably, Proposition 3.1 can be enhanced by relaxing this requirement, as detailed in [15], Section 4. This improvement hinges on extending the Gromov-weak topology to non-separable ultrametric spaces through a de Finetti representation for exchangeable coalescents, as obtained in [14].